A can of paint will cover 200ft^2. How many can of paint should be purchased in order to paint a cylinder that has a height of 15 ft and a radius of 6ft?
Answer by Cromlix(4381)

Surface area of an enclosed (top and bottom) cylinder
SA = 2*Pi*r^2 + 2*Pi*r*h
= 2*Pi*6^2 + 2*Pi*6*15
= 791.68ft^2
791.68ft^2/200 = 3.9 cans
= 4 cans of paint.
If the cylinder is open ended
SA = 2*Pi*r*h
= 2*Pi*6*15
= 565.49ft^2
565.49/200 = 2.8 cans
= 3 cans of paint
Hope this helps.
:-)